OpenStudy (anonymous):

find a power series representation for the function amd determine the radius of convergence f(x)=(x^2+x)/(1-x)^3

OpenStudy (freckles):

ok have you found the power series representation?

OpenStudy (freckles):

\[\frac{1}{(1-x)^2}=\frac{d}{dx}\frac{1}{(1-x)}\] \[\frac{1}{(1-x)^3}=\frac{1}{2} \frac{d}{dx} \frac{1}{(1-x)^2}\] so it looks like we need the power series of 1/(1-x) then we need the second derivative of that power series then multiply it by 1/2 to get the power series of 1/(1-x)^3

OpenStudy (freckles):

then when we are done multiply by the (x^2+x)

OpenStudy (anonymous):

so what does that series look like? is it just \[x^{2}+x \Sigma x^n\]

OpenStudy (freckles):

well not exactly you need to find derivative of that one series like twice

OpenStudy (freckles):

\[\frac{1}{1-x}=1+x+x^2+x^3+x^4 + \cdots \\ \frac{1}{(1-x)^2}=0+1+2x+3x^2+4x^3 + \cdots \] now you still need to differentiate one more time

OpenStudy (freckles):

i will let you do that

OpenStudy (freckles):

differentiate both sides

OpenStudy (freckles):

basically the terms in first line is x^n derivative of that is nx^{n-1} and derivative of that derivative is n(n-1)x^(n-2)

OpenStudy (freckles):

you can actually use that to write in sigma notation

OpenStudy (freckles):

@n.sorge are you still there?

OpenStudy (anonymous):

so would n=2

OpenStudy (freckles):

?

OpenStudy (anonymous):

in the sigma notation

OpenStudy (freckles):

n is n

OpenStudy (anonymous):

does it start at 0 or 2 i mean

OpenStudy (freckles):

0(0-1)x^(0-2) + 1(1-1)x^(1-2) + 2(2-1)x^(2-2) + so on...

OpenStudy (freckles):

either

OpenStudy (freckles):

those first two terms are 0

OpenStudy (anonymous):

the first two terms of the series right

OpenStudy (anonymous):

then 2+6x.....

OpenStudy (freckles):

\[\frac{1}{1-x}=1+x+x^2+x^3+x^4+ \cdots +x^{n}+\cdots \\ \text{ differentiating both sides } \\ \frac{1}{(1-x)^2}=0+1+2x+3x^2+4x^3+\cdots + nx^{n-1}+\cdots \\ \text{ differentiating both sides } \\ \frac{2}{(1-x)^3}=0+0+2+6x+12x^2+ +\cdots n(n-1)x^{n-2}+c\dots \]

OpenStudy (freckles):

but you weren't looking for 2/(1-x)^3 you were looking for 1/(1-x)^3

OpenStudy (freckles):

that c there is a type-o

OpenStudy (freckles):

ignore it

OpenStudy (freckles):

so divide both sides of that last equation by 2

OpenStudy (freckles):

\[\frac{1}{1-x}=1+x+x^2+x^3+x^4+ \cdots +x^{n}+\cdots \\ \text{ differentiating both sides } \\ \frac{1}{(1-x)^2}=0+1+2x+3x^2+4x^3+\cdots + nx^{n-1}+\cdots \\ \text{ differentiating both sides } \\ \frac{2}{(1-x)^3}=0+0+2+6x+12x^2+ \cdots +n(n-1)x^{n-2}+\cdots \]

OpenStudy (anonymous):

so 0+0+1+3x+...

OpenStudy (freckles):

\[\frac{1}{(1-x)^3}=\frac{0}{2}+\frac{0}{2}+\frac{6x}{2}+\frac{12x^2}{2}+\cdots +\frac{n(n-1)x^{n-2}}{2}+\cdots \] then finally you should know that we are looking for (x^2+x)/(1-x)^3 so take the series from above and multiply it by the (x^2+x)

OpenStudy (anonymous):

so what would that look like would you have to distribute that

OpenStudy (freckles):

n.sorge i have basically done the whole problem for you it should be pretty easy to write what I have up there in sigma notation and then take that sum and multiply it by (x^2+x)

OpenStudy (freckles):

\[\frac{1}{(1-x)^3}=\sum_{n=0}^{\infty}\frac{n(n-1)x^{n-2}}{2}\]

OpenStudy (freckles):

your function is \[\frac{x^2+x}{(1-x)^3}\] so just mutliply both sides of that equation I just wrote by (x^2+x) to get the power series for (x^2+x)/(1-x)^3